OpenStudy (anonymous):

-17π/3 Find the exact values of all six trig functions

OpenStudy (anonymous):

sin(-17pi/3)=0.866025404 cos(-17pi/3)=0.5 tan(-17pi/3)=1.73205081 cot(-17pi/3)=0.577350269 csc(-17pi/3)=1.15470054 sec(-17pi/3)=2 can I have a medal please?

OpenStudy (anonymous):

\[-17 \pi/3 = 6 \pi - 17 \pi/3 = \pi / 3\] Now you know!

OpenStudy (anonymous):

thank you, I didn't know how to reduce that!

OpenStudy (anonymous):

np